OpenStudy (anonymous):

Amelie's cat is 6 years older than her dog. The sum of their ages is 14 years. How old is Amelie's dog? years

OpenStudy (anonymous):

4 years

jhonyy9 (jhonyy9):

cat years = dog years +6 dog years =x so dog years = 4 cat years = x+6 so cat years = 4+6 =10 x+6 +x = 14 2x+6=14 2x=14-6=8 x=8/2 = 4